OpenStudy (anonymous):

how would I graph 2x -3=0

OpenStudy (shayaan_mustafa):

Hi rook :) Its a line. Do you know standard equation of a line?

OpenStudy (anonymous):

make x the subject of formula first..zen just draw a straight line at this particular value...so \[x =3/2\]

OpenStudy (anonymous):

forget the line equation and dont have my book

OpenStudy (shayaan_mustafa):

Its OK friend. I tell you :) It is\[\large y=mx+b\]

OpenStudy (shayaan_mustafa):

so write your question as a function of x i.e.\[\large f(x)=mx+b\]now according to your question it should be look like this buddy.\[\large f(x)=2x-3\]where, m=2 is a slope and b=-3 is a y-intercept.

OpenStudy (anonymous):

could you draw a sample ? I dont know what b is . M is the slope ?

OpenStudy (shayaan_mustafa):

Yeah sure I will do it for you. But we have to gather three things in order to plot a line. Until now we have gather two things one is SLOPE and other is Y-INTERCEPT. One more thing we need i.e. X-INTERCEPT.

OpenStudy (shayaan_mustafa):

Now for x-intercept we will use your given conidtion i.e. 2x-3=0 find out the value of x from here. 2x=3 x=3/2 x=1.5 so x-intercept is x=1.5

OpenStudy (shayaan_mustafa):

Now we have slope=m=2 x-intercept=1.5 y-intercept=-3 Now we have all things now we can plot our line :D

OpenStudy (shayaan_mustafa):